# Reflectionless modes as a source of Weyl nodes in multiterminal Josephson junctions

### v1

@@ Abstract
In Fig. 2(b), we show the topological phase diagram $\Lambda(t,\epsilon)$ of the junction. Two topological phase boundaries are easily observed as they correspond to straight lines following the relation $\epsilon=t$ (I) and $\epsilon=-2t$ (II). Note, that similarly as to before, this corresponds to the eigenvalues of the normal Hamiltonian to be zero, namely $E_{-}=\epsilon-t=0$ (doubly degenerate) and $E_{+}=\epsilon+2t=0$. As before, this case is associated with a vanishing self-energy $\Sigma=0$ in Eq. 15 for $\bm{\phi}=(0,\pi,0,\pi)$. It follows that $0=\det(H_{\rm eff})=-\det(H_{\rm N})^{2}$ which is zero for $E_{\pm}=0$. We find the corresponding reflection matrix where incoming and outgoing self-energies cancel, namely
$r_{145}\equiv\begin{pmatrix}{S_{\rm N}}_{1,1}&{S_{\rm N}}_{1,4}&{S_{\rm N}}_{1,5}\\ {S_{\rm N}}_{4,1}&{S_{\rm N}}_{4,4}&{S_{\rm N}}_{4,5}\\ {S_{\rm N}}_{5,1}&{S_{\rm N}}_{5,4}&{S_{\rm N}}_{5,5}\\ \end{pmatrix}.$ (16)
which results in $\det(r_{145}^{0})\propto\det(H_{\rm N})$. It follows that $\det(r_{145}^{0})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(H_{\rm eff})=0$.
There are two other boundaries (III) and (IV) in Fig. 2(b) which cross each other, leading to a more involved phase diagram. For the dashed line (III), we find that $\Sigma_{2}=0$ with $-\Sigma_{3}=\Sigma_{1}=2$ for $\bm{\phi}=(0,0,\pi,\pi)$. We find the corresponding reflection matrix
$r_{123}\equiv\begin{pmatrix}{S_{\rm N}}_{1,1}&{S_{\rm N}}_{1,2}&{S_{\rm N}}_{1,3}\\ {S_{\rm N}}_{2,1}&{S_{\rm N}}_{2,2}&{S_{\rm N}}_{2,3}\\ {S_{\rm N}}_{3,1}&{S_{\rm N}}_{3,2}&{S_{\rm N}}_{3,3}\\ \end{pmatrix}.$ (17)
with the incoming and outgoing self-energies $\Sigma_{\rm in}={\rm diag}(2,1,0)$ and $\Sigma_{\rm out}={\rm diag}(0,1,2)$, which results in the same effective contribution as for the phase dependent self-energies, namely $\Sigma_{\rm in}-\Sigma_{\rm out}=\Sigma(0,0,\pi,\pi)={\rm diag}(2,0,-2)$. With a similar calculation as in Eq. 10, we find that $\det(H_{\rm eff})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(r_{123}^{0})=0$. Additionally, as the effect of SC vanishes in the second dot, we can dress it onto the two remaining dots, resulting in an additional self-energy and effective two-dot model [77]
@@
with the effective 2-dot normal Hamiltonian $H_{\rm N}^{\rm 2dot}=\tilde{\epsilon}\sigma_{0}+\tilde{t}\sigma_{x}$ with the renormalized parameters $\tilde{\epsilon}=\epsilon-t^{2}/\epsilon$ and $\tilde{t}=t-t^{2}/\epsilon$ with the effective 2-dot self-energy $\Sigma^{\rm 2dot}=2\sigma_{z}$. In particular, upon redefining parameters, we obtain the exact same condition as before for the reflectionless mode coupling two levels, namely $\tilde{\epsilon}^{2}-\tilde{t}^{2}+4=0$, relating the three-dot RSM to one of an effective two-dot system.
The dotted line (IV) does not correspond to a RSM of the normal state scattering matrix. Yet, it can be inferred by constraints on the self-energies. Namely, consider the case when $\bm{\phi}=(0,\pm 2\pi/3,\mp 2\pi/3,0)$, resulting in the self-energies $\Sigma_{1}=e^{\pm\imath\pi/3}$, $\Sigma_{2}=-1$ and $\Sigma_{3}=e^{\mp\imath\pi/3}$ with the corresponding self-energy diagram shown in Fig. 2(c). As $|\Sigma_{1,2,3}|=1$, we can use Schur’s determinant formula to obtain the following formula [77]
@@
where we again obtain the formula for a RSM of a 2-level system with the renormalized parameters $\hat{\epsilon}=\epsilon+t/2$ and $\hat{t}=3t/2$, despite it not being possible to assign a zero-RSM to this phase boundary.
@@
As in the previous example, in Fig. 2(d) we portray the vector field $\bm{V}_{123}$ of the reflectionmatrix $r_{123}$ (see Eq. 13) for the different parameter configurations shown in the insets and portrayed in panel (b). It is seen that each parameter configuration hosts a RSM, and for the phase transition, the RSM becomes a zero-RSM at zero real energy.
Discussion.– In summary, we have shown that zero-RSMs result in topological phase boundaries of MTJJs. In particular, we derive formulas for every arising phase boundary in the two systems in this work, where most are related to zero-RSMs and thereby also related to effective two-terminal Josephson junctions tuned to phase differences of $\pi$.
However, it is not evident how these concepts generalize to non-symmetric junctions. As topology results in stability over a finite parameter range, non-trivial topology remains stable upon added asymmetry. We infer that in nearly symmetric junctions, which are close in parameter space to symmetric junctions, the topology is rooted in zero-RSMs. However, to what degree this protection is guaranteed and whether every source of non-trivial topology can always be traced back to RSMs is not clear and requires additional effort, certainly involving numerical considerations. Namely, we find phase boundaries not corresponding to RSMs, but follow very similar functional dependencies. In this context, we like to point out the similarities between our findings and former works [78, 79, 80, 81, 82, 83] that associate topological phase transitions by zero eigenvalues and resonances of the reflection matrix or zero crossings and parity transitions. Yet, it is not clear how these concepts can all be uniformly generalized to MTJJs. In conclusion, our works provides insight into the topological properties of MTJJs. We find that zero-RSMs lead to topological phase boundaries. In addition, our work provides substantial insight into model systems, which were recently used to explain measurements, which predict topology. Our results could provide a useful tool in fabricating topologically non-trivial devices, by relating topology in MTJJs to the properties of the respective normal scattering matrix.
@@
D.C.O. acknowledges support by the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation) via SFB 1432 (Project No. 425217212).

### v3

@@ Abstract
In Fig. 1(e), we show the topological phase diagram $\Lambda(t,\epsilon)$ of the junction. Two topological phase boundaries are easily observed as they correspond to straight lines following the relation $\epsilon=t$ (I) and $\epsilon=-2t$ (II). Similar to before, this corresponds to the eigenvalues of the normal Hamiltonian to be zero, namely $E_{-}=\epsilon-t=0$ and $E_{+}=\epsilon+2t=0$. Likewise, this case is associated with a vanishing self-energy $\Sigma=0$ in Eq. 12 for $\bm{\phi}=(0,\pi,0,\pi)$. It follows that $0=\det(H_{\rm eff})=-\det(H_{\rm N})^{2}$ which is zero for $E_{\pm}=0$. We find the corresponding reflection matrix $(r_{145})_{ij}=(S_{\rm N})_{ij}$ with $i,j=1,4,5$ where incoming and outgoing self-energies cancel, which results in $\det(r_{145}^{0})\propto\det(H_{\rm N})$. It follows that $\det(r_{145}^{0})=0\Leftrightarrow\det(H_{\rm N})=0\Leftrightarrow\det(H_{\rm eff})=0$.
There are two other boundaries (III) and (IV) in Fig. 1(e) which cross each other, leading to a more involved phase diagram. For the dashed line (III), it holds that $\Sigma_{2}=0$ with $\Sigma_{3}=\Sigma_{1}=2$ for $\bm{\phi}=(0,0,\pi,\pi)$. We find the corresponding reflection matrix $(r_{123})_{ij}=(S_{\rm N})_{ij}$ for $i,j=1,2,3$ with the incoming and outgoing self-energies $\Sigma_{\rm in}={\rm diag}(2,1,0)$ and $\Sigma_{\rm out}={\rm diag}(0,1,2)$, which results in the same effective contribution as for the phase dependent self-energies, namely $\Sigma_{\rm in}-\Sigma_{\rm out}=\Sigma(0,0,\pi,\pi)={\rm diag}(2,0,-2)$. With a similar calculation as in Eq. 7, we find that $\det(H_{\rm eff})=0\Leftrightarrow\det(r_{123}^{0})=0$. Additionally, as the effect of SC vanishes in the second dot, we can dress it onto the two remaining dots, resulting in an effective two-dot model [SM]
@@
with the effective 2-dot normal Hamiltonian $H_{\rm N}^{\rm 2dot}=\tilde{\epsilon}\sigma_{0}+\tilde{t}\sigma_{x}$ with the renormalized parameters $\tilde{\epsilon}=\epsilon-t^{2}/\epsilon$ and $\tilde{t}=t-t^{2}/\epsilon$ with the effective 2-dot self-energy $\Sigma^{\rm 2dot}=2\sigma_{z}$. In particular, upon redefining parameters, we obtain the exact same condition as before for the RSM coupling two levels, namely $\tilde{\epsilon}^{2}-\tilde{t}^{2}+4=0$, relating the three-dot RSM to one of an effective two-dot system.
The dotted line (IV) does not correspond to a RSM of the normal state scattering matrix. Yet, it can be inferred by constraints on the self-energies. Namely, consider the case when $\bm{\phi}=(0,\pm 2\pi/3,\mp 2\pi/3,0)$, resulting in the self-energies $\Sigma_{1}=e^{\pm\imath\pi/3}$, $\Sigma_{2}=-1$ and $\Sigma_{3}=e^{\mp\imath\pi/3}$ with the corresponding self-energy diagram shown in Fig. 1(f). As $|\Sigma_{1,2,3}|=1$, we can use Schur’s determinant formula to obtain the following formula [SM]
@@
where we again obtain a similar formula to the one of a RSM of a 2-level system with the renormalized parameters $\hat{\epsilon}=\epsilon+t/2$ and $\hat{t}=3t/2$.
@@
In Fig. 1(h) we portray the vector field $\bm{V}_{123}$ of the reflectionmatrix $r_{123}$ (see Eq. 10) for the different parameter configurations shown in the insets and portrayed in panel (b). It is seen that each parameter configuration hosts a RSM, and for the phase transition, the RSM becomes a zero-RSM at zero real energy.
We want to emphasize that these findings are also meaningful for non-symmetric junctions. As topology results in stability over a finite parameter range, non-trivial topology remains stable upon added asymmetry. We infer that in nearly symmetric junctions, which are close in parameter space to symmetric junctions, the topology is rooted in zero-RSMs.
Discussion.– In summary, we have shown that zero-RSMs result in topological phase boundaries of MTJJs. In particular, we derive formulas for every arising phase boundary in the two systems in this work, where most are related to zero-RSMs and thereby also related to effective two-terminal Josephson junctions tuned to phase differences of $\pi$. Furthermore, we have established an effective bulk-boundary correspondence for MTJJs, where we find unity transmission modes in the normal state scattering matrix on topological phase boundaries. We want to point out the similarities between our findings for MTJJs and former works [Fulga2011, Beenakker2011, Pikulin2012, Fulga2012, Beenakker2013, Beenakker2013a] that associate topological phase transitions by zero eigenvalues and resonances of the reflection matrix or zero crossings and parity transitions. Further analysis could focus on computing topological phase diagrams for every possible normal state scattering matrix, but even in the simplest case of $4\times 4$ matrices, the configuration space is very large and advanced numerical methods would be required beyond the scope of this work.
In conclusion, our works provides insight into the topological properties of MTJJs. We find that zero-RSMs of the normal state lead to topological phase boundaries. In addition, our work provides substantial insight into model systems, which were recently used to explain measurements in accordance with topology. Our results could provide a useful tool in fabricating topologically non-trivial devices, by relating topology in MTJJs to the properties of the respective normal scattering matrix.
@@
We acknowledge support by the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation) via SFB 1432 (Project No. 425217212).
blabla
I Appendix
Analysis of the normal-state scattering matrix.– In the following, we provide additional information on the normal state scattering matrix and the corresponding zero energy reflectionless scattering modes (zero-RSMs). The normal state scattering matrix of the junction can by computed via the Fisher-Lee relation [Fisher1981] from the retarded central Green’s function $g_{\rm C}=\lim\limits_{\eta\to 0}(E-H_{\rm N}+iWW^{\dagger})^{-1}$
$S_{\rm N}=1-2iW^{\dagger}g_{\rm C}W,$ (16)
with the coupling matrix $W$.
For the two-dot model in the main text, the coupling matrix is given as
$W=\sqrt{\Gamma}\begin{pmatrix}1&1&0&0\\ 0&0&1&1\end{pmatrix}$. (17)
Diagonalizing the normal scattering matrix ($D_{S}=US_{\rm N}U^{-1}$) results in
$D_{S}={\rm diag}\left(1,1,\frac{\epsilon_{+}^{*}}{{\epsilon_{+}}},\frac{\epsilon_{-}^{*}}{{\epsilon_{-}}}\right)\equiv{\rm diag}\left(1,1,d_{+},d_{-}\right)$ (18)
where $\epsilon_{\pm}=\epsilon\pm t-2\imath\Gamma$ are the eigenvalues of the effective non-hermitian normal Hamiltonian $H_{\rm N}^{\rm nh}=(\epsilon-2\imath\Gamma)\sigma_{0}+t\sigma_{x}$. To get more insight, we can express the normal scattering matrix as a function of its eigenvalues $d_{\pm}$, $S_{\rm N}=U^{-1}D_{S}(d_{+},d_{-})U$. Now, we split the system into two parts and treat the system as an effective two terminal junction. For that, consider the reflection matrices $r_{12}$ and $r_{13}$ as defined in the main text. For a zero-RSM, one eigenvalue of the reflection matrix must be zero. Thus, we compute the diagonalized forms of the respective reflection matrices which read
$D_{r_{12}}$ $={\rm diag}\left(1,\frac{1}{2}(d_{+}+d_{-})\right),$ (19)
$D_{r_{13}}$ $={\rm diag}\left(\frac{1}{2}\left(1+d_{+}\right),\frac{1}{2}\left(1+d_{-}\right)\right).$ (20)
We obtain zero-RSMs when $d_{+}=-d_{-}$ for $r_{12}$ and $d_{+,-}=-1$ for $r_{13}$. Note, that these conditions relate to the convex hull spanned by the eigenvalues of the normal scattering matrix to enclose the origin on one of its borders. Furthermore, these conditions result in conditions for the eigenvalues of the effective non-hermitian Hamiltonian, ${\rm Re}(\epsilon_{+}{\epsilon_{-}}^{*})=0$ for $r_{12}$ and ${\rm Re}(\epsilon_{\pm})=0$ for $r_{13}$ which then lead to the relations involving the eigenvalues $E_{\pm}=\epsilon\pm t$ of the normal metal Hamiltonian $H_{\rm N}$, $E_{+}E_{-}+4=0$ ($t^{2}=\epsilon^{2}+4$) and $E_{\pm}=0$ ($\epsilon=\pm t$) respectively. As discussed in the main text, RSMs correspond to a unity transmission eigenvalue. This arises from the fact that $r$ and $r^{\dagger}$ share the eigenvalue 0, and consequently does the hermitian matrix $rr^{\dagger}$. Lastly, the matrices $1-tt^{\dagger}$ and $rr^{\dagger}$ have the same spectrum [RevModPhys.69.731]. For completeness, the diagonalized form of respective transmission matrices read
$T_{12-34}$ $={\rm diag}\left(0,\frac{1}{4}\left(2-d_{+}{d_{-}}^{*}-{d_{+}}^{*}d_{-}\right)\right),$ (21)
$T_{13-24}$ $={\rm diag}\left(\frac{1}{4}\left(2-d_{+}-{d_{+}}^{*}\right),\frac{1}{4}\left(2-d_{-}-{d_{-}}^{*}\right)\right),$ (22)
and we find eigenvalues equal to 1 exactly when the same conditions as before are fulfilled. If a transmission eigenvalue is equal to 1, we can accordingly choose the phases in a way to get a zero energy solution exactly when the phases of incoming modes are zero and the ones of the outgoing modes are $\pi$. Hence, we find that we get zero energy ABSs when RSMs (or transparent modes) appear in the normal scattering region.
For the 3-dot junction the normal state scattering matrix can be computed using the Fisher-Lee relation with a coupling matrix
$W=\sqrt{\Gamma}\begin{pmatrix}1&1&0&0&0&0\\ 0&0&1&1&0&0\\ 0&0&0&0&1&1\\ \end{pmatrix}$ (23)
like in Eq. (16). Since no crossed Andreev reflections are assumed in the model, the normal state scattering region is described by a $6\times 6$ matrix. As before, the S-matrix can be diagonalized which yields
$D_{S}$ $={\rm diag}\left(1,1,1,\frac{\epsilon_{+}^{*}}{{\epsilon_{+}}},\frac{\epsilon_{-}^{*}}{{\epsilon_{-}}},\frac{\epsilon_{-}^{*}}{{\epsilon_{-}}}\right)$ (24)
$={\rm diag}\left(1,1,1,d_{+},d_{-},d_{-}\right)$ (25)
where the eigenvalues of the effective non-hermitian central Hamiltonian
$H_{\rm N}^{\rm nh}=\begin{pmatrix}\epsilon-2\imath\Gamma&t&t\\ t&\epsilon-2\imath\Gamma&t\\ t&t&\epsilon-2\imath\Gamma\end{pmatrix}$ (26)
are $D_{H_{\rm N}^{\rm nh}}={\rm diag}(\epsilon_{+},\epsilon_{-},\epsilon_{-})$ with $\epsilon_{+}=\epsilon+2t-2\imath\Gamma$ and $\epsilon_{-}=\epsilon-t-2\imath\Gamma$.
In this system, we find three zero-RSMs. The first two appear in the reflection matrix $r_{145}$, where we find
$D_{r_{145}}={\rm diag}\left(\frac{1}{2}\left(1+d_{+}\right),\frac{1}{2}\left(1+d_{-}\right),\frac{1}{2}\left(1+d_{-}\right)\right),$ (27)
in complete analogy to before. Namely, when an edge of the convex hull of the eigenvalues of the normal state scattering matrix lies on the origin, we get zero energy solutions. Another zero-RSM is found in the reflection matrix $r_{123}$, which doesn’t provide any simple geometric relation for the eigenvalues of the scattering matrix.
Details on 3-dot system.– As mentioned in the main text, for certain topological phase boundaries, the self-energy of a specific dot may be zero. This effectively decouples the dot from its environment at a zero crossing which enables us to dress this dot onto the remaining dots. Therefore, consider the phase boundary (III) in the 3-dot junction of the main text which has vanishing self-energy on the middle dot. For a zero solution of the effective Hamiltonian to appear in Eq. (1), we infer by Schur’s formula a zero energy solution of the descending Hamiltonian
$H_{\rm eff}^{\rm 2dot}=\begin{pmatrix}H_{\rm L}&t\sigma_{3}\\ t\sigma_{3}&H_{\rm R}\end{pmatrix}-\frac{t^{2}}{\varepsilon}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}\otimes\sigma_{3}.$ (28)
This corresponds to a renormalization of the dot energy and hopping parameter by $\varepsilon\to\tilde{\varepsilon}=\varepsilon-t^{2}/\varepsilon$ and $t\to\tilde{t}=t-t^{2}/\varepsilon$, resulting in an analogous treatment as for the 2-dot junction. Thus, from the relation $4\Gamma^{2}+\tilde{\varepsilon}^{2}-\tilde{t}^{2}=0$, the explicit form of the phase boundary reads
$\varepsilon=$ $\frac{-\tilde{t}^{2}+(2\tilde{t}-\tilde{\varepsilon})\tilde{\varepsilon}}{2\tilde{t}-\tilde{\varepsilon}}$, (29)
$t=$ $\sqrt{\varepsilon(\varepsilon-\tilde{\varepsilon})}$, (30)
resulting in the more complicated from as seen in the main text.
In the symmetric 3-dot model, for the phase boundary (IV), one has (up to reordering and a global sign) the situation
$H_{\rm N}=\begin{pmatrix}\varepsilon&t&t\\ t&\varepsilon&t\\ t&t&\varepsilon\end{pmatrix},\quad\Sigma=\begin{pmatrix}1&0&0\\ 0&\zeta&0\\ 0&0&\zeta^{2}\end{pmatrix}$,
with $1,\zeta,\zeta^{2}$ the roots of $z^{3}-1=0$. For a zero energy ABS to appear, we find
$\det\left(\Sigma+\frac{1}{\Gamma^{2}}H_{\rm N}\Sigma H_{\rm N}\right)=0,$ (31)
which reduces to
$\det\begin{pmatrix}a&b&b^{*}\\ b^{*}&a&b\\ b&b^{*}&a\end{pmatrix}$ $=a^{3}+2\mathrm{Re}(b^{3})-3|b|^{2}a=0$
$=C^{2}(\Gamma^{2}+(\varepsilon-t)^{2}),$
where we used the following abbreviations
$a$ $=\Gamma^{2}+\varepsilon^{2}-t^{2}=C+t(t-\varepsilon),$
$b$ $=\zeta^{2}t(t-\varepsilon),$
$C$ $=\Gamma^{2}+\varepsilon^{2}+\varepsilon t-2t^{2}=\Gamma^{2}+\left(\varepsilon+\frac{t}{2}\right)^{2}-\left(\frac{3}{2}t\right)^{2}.$
This makes it clear that the condition $C=0$ has to be fulfilled for a zero energy solution to appear. Interestingly, this could be inferred from a two dot model with renormalized symmetric levels $\hat{\varepsilon}=\varepsilon+t/2$ and hopping $\hat{t}=3t/2$ that has the same two distinct eigenvalues as the central three-dot Hamiltonian. In such a two-dot geometry, the topological phase boundary would be related to a zero-RSM of the normal scattering matrix.